liquid asset $6,80,000, Inventories $1,90,000 , prepaid expenses $10,000, working assets $2,00,000. calculate the current ratio and quick ratio


1
Expert's answer
2021-05-10T16:01:15-0400

Working Capital = Current Assets – Current Liabilities


200,000=(680,000+190,000+10,000)"-" Current Liabilities


Current Liabilities =680,000


Current Ratio = C.A : C.L


Current Ratio =680,000:680,000


Current Ratio =1:1



"Quick Ratio =\frac {Quick Assets} {Current Liabilities}"


Quick Ratio ="\frac {880,000}{680,000} =1.29times"